OpenStudy (anonymous):

Suppose a triangle has sides a, b, and c, and that a^2 + b^2 > c^2. Let θ be the measure of the angle opposite the side of length c. Which of the following must be true? Check all that apply. A. The triangle in question is a right triangle. B. cosθ > 0 C. cos θ< 0 D. θ is an acute angle.

OpenStudy (anonymous):

By using al-Kashi's theorem \[c^2=a^2+b^2-2ab \cos(\theta)\] \[a^2+b^2>c^2 \iff -2ab \cos(\theta) >0 \iff 2ab \cos(\theta) < 0 \iff \cos(\theta)<0\] This means the angle must be greater than a right angle... so an obtuse angle.